OpenStudy (anonymous):

A square piece of fabric has an area of 595 cm2. To the nearest tenth, what is the length of 1 side of the fabric? A. 21.3 cm B. 23.2 cm C. 24.4 cm D. 25.6 cm

OpenStudy (anonymous):

@ganeshie8

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{area of}\ \square \ square=(side)^2\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

I have t find the sqaure??

OpenStudy (mathmath333):

u have to find the side

OpenStudy (anonymous):

Oh okay

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} &\normalsize \text{area of}\ \square \ square=(side)^2\hspace{.33em}\\~\\ &595=(side)^2\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

Given the formula of an area of a square: \[(side)^2\] We can deduce that, \[595=(side)^2\] Solving it, just square both sides of the equation. Cancelling the square root of the right side of the equation and squaring the left side of the equation. \[\sqrt{595} = side\] 24.39 or 24.4 = side

OpenStudy (anonymous):

To find area the formula is A=l x w. So plug in each value. When I did all of them I got C. A=l x w 24.4x24.4=595.36 Rounded it to the nearest whole number it would be 595 cm^2 so the answer is c.

OpenStudy (anonymous):

C -.-

OpenStudy (anonymous):

I don't need help on this anymore ._.